# Cedegao E. Zhang1,*

### v1

@@ Introduction
Our findings demonstrate how people might make quick, probabilistic judgments about the expected value of complex novel problems—well before they learn narrower, task-specific strategies and representations that constitute genuine expertise—by integrating efficient, general-purpose computations for simulating multi-agent decision making and approximate probabilistic inferences. We see many avenues for future work.
One is the question of how people grasp the constraints and goals of novel tasks, in order to reason about them. Here we assumed that people understood the linguistic descriptions of the games we presented them with, but this is a nontrivial cognitive process in its own right. Although large language models were limited in their ability to reason about new games as people do, they have proven capable of semantic parsing and program synthesis, and we see opportunities for modeling language-based game understanding using LLMs together with probabilistic planning and decision models like our intuitive game theory models, in a neurosymbolic framework \@BBOPcitep\@BAP\@BBN(Austin et al., 2021; Chen et al., 2021; Collins et al., 2022; Wong et al., 2023)\@BBCP.
Here we also only modeled the earliest stages of playing and reasoning about novel games. It is plausible that people’s skill in these games will rapidly increase as they play more. A deeper understanding of such fast concept and skill acquisition would be highly valuable \@BBOPcitep\@BAP\@BBN(Lake et al., 2017)\@BBCP. More broadly, the study of how people quickly learn to play novel games could contribute to contemporary work on resource rationality \@BBOPcitep\@BAP\@BBN(Gershman et al., 2015; Lieder & Griffiths, 2020; Icard, 2023)\@BBCP and human efficient planning \@BBOPcitep\@BAP\@BBN(Callaway et al., 2022; Ho et al., 2022)\@BBCP. It is an open question to what extent our current model can properly be called “resource rational”, or indeed, to what extent and in which ways naive humans approximate answers to these kinds of intuitive questions using “resource rational” representations. Overall, we believe general principles of efficient use of cognitive resources can be further applied to studying games, novel problem solving, and beyond.
@@ Acknowledgments
We thank Graham Todd, Joy Hsu, Guy Davidson, Laura Schulz, Mauricio Barba, Yuka Machino, and Kartik Chandra for helpful conversations that informed this work. CEZ, LW, and JBT acknowledge support from AFOSR Grant #FA9550-19-1-0269, the MIT-IBM Watson AI Lab, MIT Quest for Intelligence, ONR Science of AI, and the Simons Center for the Social Brain. KMC acknowledges support from the Marshall Commission and the Cambridge Trust. AW acknowledges support from a Turing AI Fellowship under grant EP/V025279/1, The Alan Turing Institute, and the Leverhulme Trust via CFI.

### v2

@@ Introduction
Our findings demonstrate how people might make quick, probabilistic judgments about the expected value of complex novel problems—well before they learn narrower, task-specific strategies and representations that constitute genuine expertise—by integrating efficient, general-purpose computations for simulating multi-agent decision making and approximate probabilistic inferences. We see many avenues for future work. One is the question of how people grasp the constraints and goals of novel tasks, in order to reason about them. Here we assumed that people understood the linguistic descriptions of the games we presented them with, but this is a nontrivial cognitive process in its own right. Although large language models were limited in their ability to reason about new games as people do, they have proven capable of semantic parsing and program synthesis, and we see opportunities for modeling language-based game understanding using LLMs together with probabilistic planning and decision models like our intuitive game theory models, in a neurosymbolic framework \@BBOPcitep\@BAP\@BBN(Austin et al., 2021; Chen et al., 2021; Collins et al., 2022; Wong et al., 2023)\@BBCP.
Also, here we only modeled the earliest stages of playing and reasoning about novel games. It is plausible that people’s skill in these games will rapidly increase as they play more. A deeper understanding of such fast concept and skill acquisition would be highly valuable \@BBOPcitep\@BAP\@BBN(Lake et al., 2017)\@BBCP. More broadly, the study of how people quickly learn to play novel games could contribute to contemporary work on resource rationality \@BBOPcitep\@BAP\@BBN(Gershman et al., 2015; Lieder & Griffiths, 2020; Icard, 2023)\@BBCP and human efficient planning \@BBOPcitep\@BAP\@BBN(Callaway et al., 2022; Ho et al., 2022)\@BBCP. It is an open question to what extent our current model can properly be called “resource rational” and to what extent naive humans approximate answers to these kinds of intuitive questions using “resource rational” representations. Overall, we believe general principles of efficient use of cognitive resources can be further applied to studying games, novel problem solving, and beyond.
@@ Acknowledgments
We thank Graham Todd, Joy Hsu, Guy Davidson, Laura Schulz, Yuka Machino, Kartik Chandra, Tom Griffiths, and Ionatan Kuperwajs for helpful conversations that informed this work. CEZ, LW, and JBT acknowledge support from AFOSR Grant #FA9550-19-1-0269, the MIT-IBM Watson AI Lab, MIT Quest for Intelligence, ONR Science of AI, and the Simons Center for the Social Brain. KMC acknowledges support from the Marshall Commission and the Cambridge Trust. AW acknowledges support from a Turing AI Fellowship under grant EP/V025279/1, The Alan Turing Institute, and the Leverhulme Trust via CFI.
Appendix: Additional modeling details
Approximate Expert model details
Here we detail pseudocode for the approximate expert model (designed to generalize the 4-in-a-row model of human gameplay from \@BBOPcite\@BAP\@BBNvan Opheusden et al. (2023)\@BBCP). Its algorithm repeats 3 sub-procedures, SelectNode, ExpandNode, and Backpropagate, within the main procedure MakeMove, which ultimately makes a move by sampling from a softmax distribution.
Algorithm 1 Approximate Expert: MakeMove
$root\leftarrow\textrm{Node}(s)$
for $i=1$ to $num\_steps$ do
     $n\leftarrow\textrm{SelectNode}(root)$
     $\textrm{ExpandNode}(n)$
     $\textrm{Backpropagate}(n,root)$
end for
return $\textrm{Sample}(\textrm{softmax}(\{c:c.val\textrm{ for }c\in root.children\})$
The function Node is the constructor for the nodes of the tree search. In our simulations, we repeat the procedure in MakeMove for $num\_steps=636$ iterations, drawing on results from an empirical parameter fit in \@BBOPcite\@BAP\@BBNvan Opheusden et al. (2023)\@BBCP which found that expert human players (playing a specific 4-in-a-row variant) were well modeled by a heuristic search procedure whose number of iterations is geometrically distributed with a mean stopping parameter of 1/636. Our model does not use a stochastic stopping parameter and just runs deterministically to $num\_steps=636$ iterations for all simulations.
Algorithm 2 Approximate Expert: $\textrm{SelectNode}(n)$
$root\leftarrow n$
while $n.children\neq\emptyset$ do
     if $n.player\_type=root.player\_type$ then
         $n=\textrm{argmax}_{c\in n.children}c.val$
     else
         $n=\textrm{argmin}_{c\in n.children}c.val$
     end if
end while
return $n$
Here player_type $\in\{X,O\}$.
Algorithm 3 Approximate Expert: $\textrm{ExpandNode}(n)$
$s\leftarrow n.state$
for all $m\in\textrm{LegalMoves}(s)$ do
     $b\leftarrow\textrm{PlacePiece}(s,m,n.player\_type)$
     $i\leftarrow n.move\_number$
     $n.children.append(\textrm{Node}(b,i+1,V^{\textrm{AE}}(b))$
end for
Here move_number is the index of the current move over the entire game (to account for variants which specify that a given player goes twice as their opening move).
Within the procedure ExpandNode, the function $V^{\textrm{AE}}$ is a modified version of our heuristic function $V$ defined in equation 1 that depends only on the board state $b$ and is symmetric for both players. To set a frame of reference, suppose player $X$ is evaluating the utility of board state $b$. The heuristic $V^{\textrm{AE}}(b)$ is evaluated as $V_{X}^{\textrm{AE}}(b)-V_{O}^{\textrm{AE}}(b)$ where
$V^{\textrm{AE}}_{a}(b)=\sum_{p\in a\textrm{ pieces}}V(p,b).$
for $a\in\{X,O\}$. The summation is over all pieces $p$ on the board that $a$ has placed. When $O$ acts, the heuristic is symmetrically $V^{\textrm{AE}}(b)=V^{\textrm{AE}}_{O}(b)-V^{\textrm{AE}}_{X}(b)$. We make this modification so that $V^{\textrm{AE}}$ is comparable across different branches of the search tree and so that we can efficiently alternate between the two players in the tree.
Algorithm 4 Approximate Expert: $\textrm{Backpropagate}(n,root)$
if $n.player\_type=root.player\_type$ then
     $n.val\leftarrow\max_{c\in n.children}c.val$
else
     $n.val\leftarrow\min_{c\in n.children}c.val$
end if
MCTS model details
Similar to the approximate expert model, MCTS repeats 4 sub-procedures—SelectNode, ExpandNode, DepthCharge and Backpropagate, but the SelectNode, ExpandNode, and Backpropagate sub-procedures are different from their corresponding counterparts in the approximate expert model. We run MCTS for 10,000 steps, which we find balances a close approximation of the optimal for many games while keeping the runtime of any single game trial to less than 2 days.
Algorithm 5 MCTS: MakeMove
$root\leftarrow\textrm{Node}(s)$
for $i=1$ to $num\_steps$ do
     $n\leftarrow\textrm{SelectNode}(root)$
     $\textrm{ExpandNode}(n)$
     $game\_outcome=\textrm{DepthCharge}(n,root)$
     $\textrm{Backpropagate}(n,root,game\_outcome)$
end for
return $\textrm{argmax}_{c\in root.children}c.val$
Algorithm 6 MCTS: $\textrm{SelectNode}(n)$
while $n.children\neq\emptyset$ do
     $n=\textrm{argmax}_{c\in n.children}c.UCB$
end while
return $n$
To balance exploring and exploiting in the MCTS SelectNode sub-procedure, we use a standard upper confidence bound (UCB) \@BBOPcitep\@BAP\@BBN(Browne et al., 2012)\@BBCP.
$c.UCB=$ $\textrm{argmax}_{c\in root.children}\big{(}\frac{c.val}{c.visits}$
$+\sqrt{2\frac{\log(c.parent.visits)}{c.visits}}\big{)}.$
Algorithm 7 MCTS: $\textrm{ExpandNode}(n,root)$
$s\leftarrow n.state$
for all $m\in\textrm{LegalMoves}(s)$ do
     $b\leftarrow\textrm{PlacePiece}(s,m,n.player\_type)$
     $i\leftarrow n.move\_number$
     $n.children.append(\textrm{Node}(b,i+1))$
end for
Algorithm 8 MCTS: $\textrm{Backpropagate}(n,root,game\_outcome)$
$n.visits\leftarrow n.visits+1$
if n.parent then
     if n.parent.player_type == root.player_type then
         $n.val\leftarrow n.val+game\_outcome$
     else
         $n.val\leftarrow n.val+(1-game\_outcome)$
     end if
     $\textrm{Backpropagate}(n.parent,root,game\_outcome)$
end if
Algorithm 9 MCTS: $\textrm{DepthCharge}(n)$
$s\leftarrow n.state$
$i\leftarrow n.move\_number$
while not $\textrm{HasTerminated}(s)$ do
     $player\_type\leftarrow n.player\_sequence[i]$
     $m\leftarrow\textrm{RandMove}(s)$
     $i\leftarrow i+1$
     $s\leftarrow\textrm{PlacePiece}(s,m,player\_type)$
end while
if $\textrm{IsWin}(s,root.player\_type;game\_rules)$ then
     return 1
else if $\textrm{IsDraw}(s;game\_rules)$ then
     return 1/2
else
     return 0
end if
Finally, as noted above, our implementations of the approximate expert and MCTS algorithms always parameterize simulation with a formal specification of the rules of any given game (e.g., restrictions on whether a given player can win only in certain directions on the board). We implement $V$ and DepthCharge as described in the main text so that intermediate value functions and win conditions are calculated with respect to the specific game dynamics and win conditions for each game variant. MakeMove similarly takes in the current move_number move index to account for games in which player order is not strictly alternating, such as games in which the first player goes twice as their opening move.